OpenStudy (king):

: Help..............In a school a certain number of boys and girls attend a certain examination. The mean mark of the boys was 40 and the mean mark of the girls was 60.Mean of all students is 48.Find the number of boys and girls in a class of 100.

OpenStudy (wasiqss):

(40x+60y)/100=48 x+y=100

OpenStudy (wasiqss):

solve two simoultaneous eq

OpenStudy (wasiqss):

boys r 60, galx 100

OpenStudy (king):

not possible sorry wrong answer x+y needs to be 100

OpenStudy (wasiqss):

sorry i mean to say galx 40

OpenStudy (wasiqss):

typed wrong

OpenStudy (king):

oh ok but how did u get first two equations....xplain

OpenStudy (king):

sry only first equation

OpenStudy (wasiqss):

listen what will be mean of marks? total marks divided by candidate

OpenStudy (king):

and what is x and what is y?

OpenStudy (wasiqss):

x is boys , y is number of galx

OpenStudy (king):

yeah but how 40 * x and 60 * y?

OpenStudy (wasiqss):

so the total marks is number of boys into their marks+number of galx nto their mean marks then divided by total number of students that s 100

OpenStudy (wasiqss):

ok

OpenStudy (king):

why is 40 multiplied by x and 60 multiplied into y?

OpenStudy (wasiqss):

40 is the mean marks of boys and x is the total number of boys in class and y is total number of gals ad 60 is the mean marks of gal

OpenStudy (king):

yeah so why do we multiply?

OpenStudy (king):

why do we multiply?

OpenStudy (king):

why do we multiply?

OpenStudy (king):

anwar help?

OpenStudy (anonymous):

OK. Let \(n\) and \(m\) be the number if boys and girls respectively, then \(40 n\) is the total points that boys scored, and \(60m\) is the total of the girls. We can then write \(\frac{40n+60m}{m+n}=48\). We also have \(m+n=100\). Solve them to find m and n.

OpenStudy (wasiqss):

we multiply to get the total number of marks like suppose there are 3 students in class and their mean mark is 90 so how we will get the marks obtained by class in total as in we want to add the marks of all the student

OpenStudy (anonymous):

If you solve them, you will get \(n=70 \text{ boys, and } m=30 \text{ girls}\).

OpenStudy (king):

no anwar u are wrong

OpenStudy (king):

wasiq why will we multiply if 3 students are there then mmean is x1+x2+x3/3=90

OpenStudy (anonymous):

Oh sorry, I solved them for \(46\), not \(48\). Wait a sec.

OpenStudy (anonymous):

60 boys and 40 girls :)

OpenStudy (king):

yea anwser is rite but i dont understand why we multiply...

OpenStudy (asnaseer):

the mean of the these 3 numbers: 12,14,16 is calculated as:\[\frac{12+14+16}{3}=\frac{42}{3}=14\]the total of those 3 numbers is: 12+14+16=42, which is the same as the mean times the number of numbers you have: 14*3=42

OpenStudy (wasiqss):

king u r dumb, sorry to say bt tru

OpenStudy (asnaseer):

so, if you are given the mean and the number of items, you can always work out the total number of items as (mean)*(number of items)

OpenStudy (anonymous):

Let me explain it to you. Assume we have 5 students, and their mean is 50. We can write that as \(\large{\frac{x_1+x_2+x_3+x_4+x_5}{5}=50}\). Now if we want to find the sum of the points that all the 5 student scored, we just multiply both sides by \(5\) to get \(\large {x_1+x_2+x_3+x_4+x_5=5\times 50}\). Makes sense?

OpenStudy (asnaseer):

sorry "the total of all the items"

OpenStudy (king):

oh ok thnx anwar and asnaseer thnk u very much

OpenStudy (asnaseer):

no problem - always glad to help